OpenStudy (anonymous):

Rational Expression: (x-1)/(x²-1) + (2)/(5x+5) These are fractions !

OpenStudy (anonymous):

7/(5x+5)

OpenStudy (anonymous):

\[\frac{x-1}{x^2-1}=\frac{x-1}{(x+1)(x-1)}=\frac{1}{x+1}\]Also,\[\frac{2}{5x+5}=\frac{2}{5(x+1)}\]So you have\[\frac{x-1}{x^2-1}+\frac{2}{5x+5}=\frac{5}{5}\times \frac{1}{x+1}+\frac{2}{5(x+1)}=\frac{7}{5(x+1)}\]

OpenStudy (anonymous):

Thanks, Radar :)

OpenStudy (anonymous):

Awww... I had the same answer, just simplified. Lulz.

OpenStudy (anonymous):

Weird, that answer wasn't in my answer key. It's ok. Thank for helping though :)

OpenStudy (anonymous):

np